OpenStudy (anonymous):

A gas has a volume of 590 ML at a temperature of -55.0°C what volume will the gas occupy at 30.0° C I don't understand this at all, please help.

OpenStudy (matt6288):

V1T1 = V2T2 (590 mL)(-55.0C) / (30.0C) = 1081.67 mL 273 + (-55) = 218 K 590/218 = V2/303 V2 = 590 (303/218) = 820 ml

OpenStudy (matt6288):

Use V1/T1 = V2/T2 assuming constant pressure change temps to Kelvin